OpenStudy (anonymous):

if f(x)=6x-4 g(x)=x-2 how do i find (f of g)(x)

OpenStudy (anonymous):

(f of g)(x)=f(g(x))=f(x-2) now just put x-2 instead of x in f's equation

OpenStudy (anonymous):

what will u get?

OpenStudy (anonymous):

6x(x-2)-4?? is that my equation

OpenStudy (unklerhaukus):

almost /

OpenStudy (unklerhaukus):

\[f(x)=6(x)-4\] \[f(a+b)=6(a+b)-4\]

OpenStudy (anonymous):

6(x-2)-4??

OpenStudy (unklerhaukus):

yeah! now you can simplify a little bit and your done

OpenStudy (anonymous):

6x-8

OpenStudy (unklerhaukus):

not quite

OpenStudy (anonymous):

after 6(x-2)-4 should i get 6x-12-4..

OpenStudy (anonymous):

6x=16?

OpenStudy (unklerhaukus):

\[f\circ g(x)=6x-16\] is right

OpenStudy (anonymous):

yay okay thanksss

OpenStudy (anonymous):

im going to close this question, and im going to ask another one..*hince* i want you to help me again :) if you can ha

OpenStudy (unklerhaukus):

ill see what i can do